an express train travels 600 mi in the same amount of time it takes a freight to travel 360 mi. the rate of the express train is 20 mph faster than that of the freight train. find the rate of each train.
Answer by mananth(12270)

Freight train x ---360 miles
Express x+20--- 600 miles
360/x=600/(x+20)
Cross multiply
360(x+20 )=600* x
360x+7200=600x
600x-360 x=7200
240x=7200
/240
x=30 mph Freight train
50 mph Express
m.ananth@hotmail.ca